OpenStudy (anonymous):

Tania bought a glass pyramid from a gift store. The pyramid, with a height of 4 cm and a square base of side 6 cm, was packed in a box shaped as a cube of side 8 cm. The empty space in the box was packed with paper tissue. How many cubic centimeters of the box were stuffed with paper tissue? Answer 192 cubic centimeters 368 cubic centimeters 464 cubic centimeters 480 cubic centimeters

OpenStudy (anonymous):

Volume of cubical box minus Volume of square pyramid. Volume of cubical box = 8^3 = 512 cubic centimeters. Volume of square pyramid = 1/3 * base^2 * h = 1/3 * 6^2 * 4 = 48 cubic centimeters. So, the answer is: 512 - 48 = 464 cubic centimeters.